Edexcel IGCSE Math A Higher 6.3 Probability Questions
Practise probability language, scales, event lists, Venn diagrams and calculations using fractions, decimals and complements.
- Syllabus
- First assessment 2018
- Course
- Math A 4MA1
- Level
- Higher
Practise probability language, scales, event lists, Venn diagrams and calculations using fractions, decimals and complements.
A biased spinner has three sections each of a different colour.
The table shows the probability that, when the spinner is spun once, it will land on blue or on orange or on white.
Work out the value of x
The spinner is spun 250 times.
Remaining probability:
1-0.58=0.42
Since the other probabilities are 2x and x:
0.58+2x+x=13x=0.42
Final answer:
x=0.14
Equivalent answers accepted, for example 14% or 507.
Work out an estimate for the number of times the spinner will land on blue.
Estimate:
250×0.58=145
Final answer:
145
M1 for 250×0.58 or equivalent.
Grace is going to spin the arrow on the spinner once.
The table below gives the probabilities that the spinner will land on red or on blue or on green.
The probability that the spinner will land on orange is 3 times the probability that the spinner will land on pink.
Work out the probability that the spinner will land on orange.
Working:
1-(0.2+0.12+0.08)=0.60.6×43=0.45
Final answer:
0.45
M1 for a correct calculation or equation for the remaining probability. M1 for finding 43 of the remaining probability. A1 for 0.45, 209, 45%, or 10.45.
Grace spins the arrow on the spinner 150 times.
Work out an estimate for the number of times the spinner lands on blue.
Working:
0.12×150=18
Final answer:
18
M1 for a correct calculation for the number of times the spinner lands on blue. A1 for 18. An answer of 15018 scores M1A0 because it is a probability, not a number of times.
100 farmers are asked if they have goats (G), sheep (S) or chickens (C) on their farms.
Of these farmers
31 have sheep
53 have chickens
6 have goats, sheep and chickens
11 have sheep and goats
17 have sheep and chickens
18 have goats and chickens
20 do not have any goats, sheep or chickens
Find the probability that this farmer also has goats.
Final answer:
5318
Accept 0.33… or 33.96…%. Follow through from the candidate's Venn diagram.
A biased spinner has three sections each of a different colour.
The table shows the probability that, when the spinner is spun once, it will land on blue or on orange or on white.
Work out the value of x
The spinner is spun 250 times.
Remaining probability:
1-0.58=0.42
Since the other probabilities are 2x and x:
0.58+2x+x=13x=0.42
Final answer:
x=0.14
Equivalent answers accepted, for example 14% or 507.
Grace is going to spin the arrow on the spinner once.
The table below gives the probabilities that the spinner will land on red or on blue or on green.
The probability that the spinner will land on orange is 3 times the probability that the spinner will land on pink.
Work out the probability that the spinner will land on orange.
Working:
1-(0.2+0.12+0.08)=0.60.6×43=0.45
Final answer:
0.45
M1 for a correct calculation or equation for the remaining probability. M1 for finding 43 of the remaining probability. A1 for 0.45, 209, 45%, or 10.45.
The table shows information about the weights, in kilograms, of 40 babies.
Find the probability that this baby has a weight of more than 5 kg .
402+401=403
Final answer:
403
Equivalent:
0.075
Ciara throws four fair six-sided dice.
The faces of each dice are labelled with the numbers 1,2,3,4,5,6
Work out the probability that at least one of the dice lands on an even number.
Probability no dice lands on an even number:
(21)4=161
Therefore:
1−(21)4=1615
Final answer:
1615
Equivalent decimals accepted, for example 0.9375.
Here is a biased spinner.
When the spinner is spun once, the probabilities that it lands on red or on yellow or on green are given in the table.
Work out the probability that the spinner lands on red or on yellow.
0.45
1
B1 oe eg 209,10045,45%
Grace is going to spin the arrow on the spinner once.
The table below gives the probabilities that the spinner will land on red or on blue or on green.
The probability that the spinner will land on orange is 3 times the probability that the spinner will land on pink.
Grace spins the arrow on the spinner 150 times.
Work out an estimate for the number of times the spinner lands on blue.
Working:
0.12×150=18
Final answer:
18
M1 for a correct calculation for the number of times the spinner lands on blue. A1 for 18. An answer of 15018 scores M1A0 because it is a probability, not a number of times.
Emilie takes part in two races.
The probability that she wins the first race is 0.7
The probability that she wins the second race is 0.4
The outcomes of the two races are independent.
Complete the probability tree diagram.
First race
Second race
First race missing probability:
0.3
Second race branch probabilities:
0.6,0.4,0.6
B1 for the first race branch, B1 for the second race branches.
Emilie takes part in two races.
The probability that she wins the first race is 0.7
The probability that she wins the second race is 0.4
The outcomes of the two races are independent.
Work out the probability that Emilie wins exactly one of the two races.
Emilie is going to take part in a third race.
If she wins both of the first two races, the probability that she will win the third race is 0.6 If she wins exactly one of the first two races, the probability that she will win the third race is 0.3
Exactly one win:
0.7×0.6+0.3×0.4=0.42+0.12=0.54
Final answer:
0.54
Equivalent answer 5027 accepted.
Here are 9 cards. Each card has either a number on it or a letter on it.
Tomas is playing a game.
Tomas takes at random one of the cards and keeps it.
Tomas then takes at random another card and keeps it.
Work out the probability that each of the two cards has a number on it.
92×81=722=361
Final answer:
361
Equivalent decimal to at least 2 significant figures accepted.
Work out the probability that there will be one card with a number on it and one card with a letter on it.
One number and one letter:
92×87+97×82=7214+7214=187
Final answer:
187
Equivalent decimal to at least 2 significant figures accepted.
Emilie takes part in two races.
The probability that she wins the first race is 0.7
The probability that she wins the second race is 0.4
The outcomes of the two races are independent.
Work out the probability that Emilie will win exactly two of the three races.
Exactly two wins in three races:
0.7×0.4×(1−0.6)=0.1120.54×0.3=0.1620.112+0.162=0.274
Final answer:
0.274
Equivalent answer 500137 accepted.